Assume you want to receive Tk. 25,000 per year for next five (5) years to cover 

your educational expenses. To get your desired cash flow how much you need 

to deposit today in your local bank which pays 11% profit per year?


1
Expert's answer
2021-03-09T02:17:07-0500

"Present \,value=PMT\times\frac{1-(1
+r)^{-n}}{r}"


where; PMT=25,000, r=11% and n=5


"Present \,value=25,000\times\frac{1-(1
+0.11)^{-5}}{0.11}=92,397"


Amount to invest now= Tk. 97,397